# NRSurNN3dq4: A Deep Learning Powered Numerical Relativity Surrogate for Binary Black Hole Waveforms

### v1

@@ Abstract
Gravitational wave approximants are widely used tools in gravitational-wave astronomy. They allow for dense coverage of the parameter space of binary black hole (BBH) mergers for purposes of parameter inference, or, more generally, match filtering tasks, while avoiding the computationally expensive full evolution of numerical relativity simulations. However, this comes at a slight cost in terms of accuracy when compared to numerical relativity waveforms, depending on the approach. One way to minimize this is by constructing so-called surrogate models which, instead of using approximate physics or phenomenological formulae, rather interpolate within the space of numerical relativity waveforms. In this work, we introduce NRSurNN3dq4, a surrogate model for non-precessing BBH merger waveforms powered by neural networks. By relying on the power of deep learning, this approximant is remarkably fast and competitively accurate, as it can generate millions of waveforms in a tenth of a second, while mismatches with numerical relativity waveforms are restrained below $10^{-3}$. We implement this approximant within the bilby framework for gravitational-wave parameter inference, and show that it it is suitable for parameter estimation tasks.
preprint: APS/123-QED
@@ I Introduction
Gravitational-wave (GW) astronomy is a computationally intensive field. Within the LIGO-Virgo-KAGRA (LVK) collaboration aligo_2015 ; avirgo_2015 ; kagra_2019 , a number of detection pipelines (GstLAL sachdev2019gstlal , MBTA Aubin_2021 , PyCBC Dal_Canton_2021 , SPIIR chu2021spiir , and cWB Klimenko_2016 ) are constantly monitoring detector data in order to inform collaboration members of potential GW signatures, as well as facilitate public alerts for the broader astrophysical community. If a GW trigger is detected by the pipelines of the LVK collaboration, a further analysis is often desired for parameter estimation. In the case of binary black hole (BBH) events, this is done by maximizing some likelihood function based on a similarity metric between the detector data and simulated GW waveforms, typically a noise-weighted inner product, which should reach a maximum if the original wave is exactly the same as some produced simulation Maggiore_book . This means one must have at hand some process which, given some physical parameters as input, will produce a physically accurate waveform, while being fast enough to allow an exploration of the GW parameter space. Waveforms from full numerical relativity simulations, which can take weeks to complete, are not appropriate for this task. Instead, one typically uses approximants, computational routines which produce fast approximations of what a gravitational waveform should look like given some physical parameters. As they are ultimately tested against (and often calibrated according to) numerical relativity waveform catalogues (such as the ones produced by the SXS collaboration SXS ), approximants tend to be highly accurate and indeed have allowed for the analysis of all GW signals detected to date GWTC-1 ; GWTC-2 ; GWTC-2.1 ; GWTC-3 .
@@
Despite this remarkable success, there is still room for improvement in the field. Parameter estimation is typically done through Bayesian inference based on dynamic nested sampling bilby_paper . This process can be parallelized in order to speed up inference times (note, however, that this speedup is not linear with the number of cores dynesty_scaling ). This is usually done by sending Bayesian inference jobs to high-performance computing (HPC) clusters, making use of large numbers of CPU cores to obtain the desired parallelization gains. However, even for these implementations the computational bottleneck can be found in the calculation of the likelihood, in particular the generation of the waveform.
@@
Next generation GW detectors, namely Einstein Telescope and Cosmic Explorer, are expected to achieve detection rates of 100,000 BBH events per year ET_rates ; CE_rates . In order to process these events efficiently with Bayesian methods, the waveform generation process must be made faster, while still allowing for an accurate parameter estimation. In order to address this issue, applications of deep learning methods to steps of this process have been proposed. Approaches based on likelihood-free inference avoid the waveform generation step altogether by replacing Bayesian parameter estimation with neural posterior estimation dingo_paper , while others strive for a speed-up through more efficient sampling using normalizing flows nessai ; nessai_sw .
In this paper, we instead tackle the problem through the lens of speeding up waveform generation. Existing literature has shown machine learning to be an effective method for waveform generation, in particular to provide reduced order models of existing approximants. These typically involve some sort of low-dimensional representation of the waveform, such as singular value decomposition (SVD), principal component analysis (PCA) or autencoder bases ANN_Sur_FD ; SVD_meshfree ; Nousi_spiral ; schmidt_ml_gw_bbh ; Luna2024 ; fontbute2024 ; grimbergen2024generatinghigherordermodes . In this work, we expand on previous work by producing a surrogate model trained directly on numerical relativity data. We introduce NRSurNN3dq4111The NRSurNN3dq4 model name indicates its development as a neural network-powered surrogate (NN) trained for non-precessing systems (3d, as the parameter space is fully defined by 3 parameters) with mass ratios up to 4 (q4), a fast and accurate surrogate to the dominant (2,2) mode of numerical-relativity-produced GWs leveraging the GPU-based nature of modern machine learning. By training a small neural network to predict the coefficients of a reduced PCA basis of approximant waveforms, and then fine-tuning on numerical relativity waveforms, we obtain a model which is capable of generating millions of accurate waveforms on a consumer-level GPU in a matter of milliseconds.
@@
The paper is organized as follows: in Section II, we introduce the datasets used in this work, including the approximant and numerical relativity waveforms. Section III describes our methodology, detailing the dimensionality reduction approach and neural network architecture used for training. The results of our model, including accuracy, speed, and parameter estimation capabilities, are presented in Section IV. Finally, we summarize our findings and discuss potential future directions in Section V.
@@ II Datasets
In the development of deep learning methods, data availability often poses a significant constraint. This proves particularly problematic in the field of numerical relativity, where the high costs associated with simulations result in a scarcity of data. For instance, by applying constraints of an effective precessing spin $\chi_{p}<0.05$ and a mass ratio $q<6$ to the SXS BBH data SXS , we obtain only $381$ samples of numerical relativity simulations. Given that deep learning frameworks typicaly need very large datasets in order to properly interpolate between the points of the input and output parameter spaces, current numerical catalogues are inadequate to fully explore them. To compensate for this, we propose to generate a substantial GW dataset using an existing approximant to train a model which can capture the fundamental characteristics of BBH GW signals, and subsequently fine-tune this model using numerical relativity data.
II.1 Approximant dataset
@@
With the aid of the gwsurrogate package for python gwsurrogate_ref , we make use of the NRSur7dq4 surrogate model NRSur7dq8ref in order to generate 1,024,000 waveforms. We restrict ourselves to aligned-spin BBH mergers, in order to keep complexity manageable, with mass ratios sampled uniformly from the $[1,6]$ interval [^2] Note that the approximant was trained in the $[1,4]$ interval, but the creators support generation up to $q=6$, claiming acceptable accuracy. As such it makes sense to train our model up to this value as well., while the aligned (dimensionless) component spins $\chi_{1},\chi_{2}$ are sampled uniformly from the $[-0.99,0.99]$ interval. We note that while uniform sampling does not necessarily reflect population-informed priors, it avoids biasing towards any particular parameter combination. The waveforms are generated in dimensionless (simulation) units, with amplitude and time scaled by the solar mass, with a sample rate of $2~{}\mathrm{M}_{\odot}^{-1}$. The merger time is estimated as the time which maximizes the amplitude of the quadrupolar $(l,|m|)=(2,2)$ mode, and the waveforms are trimmed to a duration of $4096~{}\mathrm{M}_{\odot}$, lining up the merger $100~{}\mathrm{M}_{\odot}$ before the end of the window. We further simplify our problem by setting the initial phase of the (2,2) mode in all the waveforms to zero. The (2,2) mode of the waveform is then saved to an HDF5 file, along with the physical parameters for the generation.
@@ II.2 Numerical Relativity waveforms
For setting up the numerical relativity dataset, we make use of the publicly available BBH simulation data from the SXS collaboration SXS , consisting of 2018 simulations. The (2,2) waveforms are extracted using the SXS python package sxs_package_2024 . To match the data we produced with the approximant, we must do a number of checks. First, we take into account only the waveforms which have a duration of at least $4096~{}\mathrm{M}_{\odot}$. Then, we select the non-precessing waveforms by filtering effective precessing spins $\chi_{p}<0.05$ eff_prec_spin , with mass ratios $q<6$, so that we are in line with the approximant data described previously. This process leaves us with only 381 waveforms. These waveforms are interpolated using cubic splines to the same sample rate as the approximant dataset, and trimmed to $4096~{}\mathrm{M}_{\odot}$.
@@ III.1 Dimensionality reduction
When dealing with numerical data, gravitational waveform modes are essentially arrays of $M$-length complex numbers. They are then, formally speaking, points in $\mathbb{C}^{M}$. Taking $N=2M$, we can reframe this as points in $\mathbb{R}^{N}$. However, the high correlation between adjacent points in the waveform, as well as restrictions to the shape of the waveforms (that is, wavelike signals with a chirp and ringdown-like features) means that gravitational waveforms occupy a very restricted subspace of $\mathbb{R}^{N}$. On the other hand, neural networks with $N$ outputs will output to some generic subspace of $\mathbb{R}^{N}$, the size of which will depend on the number of layers and neurons per layer (that is, its expressivity). Therefore, we need the network to learn not only the shape of the subspace, as well as how this space is connected. Furthermore, the average Euclidean distance between two points in a cube of side $l$ in $\mathbb{R}^{N}$ scales as $O(\sqrt{N})$ avg_dist_ndim_ref .
Figure 2: Mean mismatch values as a function of the number of principal components kept for the phase.
@@
This means that, in general, high dimensional data will be more sparse than lower dimensional data. As such, a neural network will have to interpolate over longer distances in such a case, which may lead to problems with convergence and generalisation curse_of_dim . Therefore, the projection of $N$-dimensional data to some lower-dimensional space should make the data more dense and thus facilitate the ease of convergence of neural networks training on that data. This is the concept of dimensionality reduction, and it is widely used across applications of machine learning DimRed_grapes ; DimRed_phys ; DimRed_wav . There are many methods for dimensionality reduction, both linear and non-linear. Our use case involves training a model in a lower-dimensional representation space of gravitational waves, and then performing an inverse transform from that space back to the typical waveform representation. This requires that any method we choose must have an accurate inverse transform. While linear methods typically have an easily calculable inverse, non-linear methods very often do not have a tractable inverse, and they must be fitted to the data in a similar way as the forward transform, such as in autoencoders or UMAP ae_ref ; umap . In this work, we will focus on principal component analysis (PCA), one of the most successful linear dimensionality reduction methods PCA_ref ; pca_rev .
@@
PCA works by calculating a linear transform $\mathcal{L}:X~{}\rightarrow(X-\mu)V$, where $X$ is some input $N\times M$ matrix (interpreted in our case as being a collection of $N$ waveforms of length $M$), $\mu$ is the average value at each timestep, and $V$ is the $m\times n$ matrix of eigenvectors of the covariance matrix $\Sigma=\frac{1}{n}(X-\mu)^{T}(X-\mu)$. The crucial point here is that the centering of $X$ means that the eigenvalues of $\Sigma$ represent the explained variance of the corresponding eigenvector. As such, if we consider only the eigenvalue-sorted subset of the top-valued $k$ eigenvectors, $V_{k}$ (an $m\times k$ matrix), we can approximately reconstruct the original dataset as
$X^{\prime}=[(X-\mu)V_{k}]V^{T}_{k}+\mu\,.$ (1)
@@
This will have some error $E$ associated to it, which will depend on the amount of components truncated (that is, $E\rightarrow 0$ as $k~{}\rightarrow n$).
@@
In our case, we will separate the complex waveform arrays $h=h_{+}-ih_{\times}=Ae^{i\phi}$ into 2-channel arrays, separating the amplitude $A$ and (unwrapped) phase $\phi$ quantities, as represented in Figure 1. This allows for a simpler, numerically non-oscillating representation, and also allows us to perform PCA separately on each quantity. Therefore, we can tune the amount of precision required in the reconstruction of each quantity. Specifically, we know that minimizing the error in the phase reconstruction is paramount for a waveform generator which is intended to be used in GW astronomy. Figure 2 shows the average mismatch values (see definition below) as a function of the number of principal components kept for the phase. Going forward, our chosen representation will keep 30 components for the amplitude and 70 components for the phase, thus reducing the 2048-dimensional complex representation of waveforms (which is equivalent to a 4096-dimensional representation) to a 100-dimensional one.

### v3

# A Deep Learning Powered Numerical Relativity Surrogate for Binary Black Hole Waveforms

@@ Abstract
Gravitational-wave (GW) approximants are essential for gravitational-wave astronomy, allowing the coverage of the binary black hole parameter space for inference or match filtering without costly numerical relativity (NR) simulations, but generally trading some accuracy for computational efficiency. To reduce this trade-off, NR surrogate models can be constructed using interpolation within NR waveform space. We present a 2-stage training approach for neural network-based NR surrogate models. We initially train four models on waveforms generated from four different GW approximants, and then fine-tune these models on NR data. We show that despite the median mismatches of the pre-trained models with NR ranging over two orders of magnitude, the fine-tuned models all reach median mismatches of order $10^{-5}$, on par with top-performing NR surrogates. The dual-stage artificial neural surrogate (DANSur3dq8) models also offer rapid waveform generation, with millions of waveforms being generated in under 20ms on a GPU. Implemented in the bilby framework, we show DANSur3dq8  can be used for parameter estimation tasks.
@@ I Introduction
Gravitational-wave (GW) astronomy is a computationally intensive field. Within the LIGO-Virgo-KAGRA (LVK) collaboration [1, 2, 3], a number of detection pipelines (GstLAL [4], MBTA [5], PyCBC [6], SPIIR [7], and cWB [8]) are constantly monitoring detector data in order to inform collaboration members of potential GW signatures, as well as facilitate public alerts for the broader astrophysical community. If a GW trigger is detected by the pipelines of the LVK collaboration, a further analysis is often desired for parameter estimation. In the case of binary black hole (BBH) events, this is done by maximizing some likelihood function based on a similarity metric between the detector data and simulated GW waveforms, typically a noise-weighted inner product, which should reach a maximum if the original wave is exactly the same as some produced simulation [9]. This means one must have at hand some process which, given some physical parameters as input, will produce a physically accurate waveform, while being fast enough to allow an exploration of the GW parameter space. Waveforms from full numerical relativity simulations, which can take weeks to complete, are not appropriate for this task. Instead, one typically uses approximants, computational routines which produce fast approximations of what a gravitational waveform should look like given some physical parameters. As they are ultimately tested against (and often calibrated according to) numerical relativity waveform catalogues (such as the ones produced by the SXS collaboration [10]), approximants tend to be highly accurate and indeed have allowed for the analysis of all GW signals detected to date [11, 12, 13, 14].
@@
Despite this remarkable success, there is still room for improvement in the field. Parameter estimation methods relying on repeated waveform evaluations, such as bilby’s Bayesian inference with dynamic nested sampling [15], benefit greatly from fast waveform generation. This process can be parallelized in order to speed up inference times (note, however, that this speedup is not linear with the number of cores [16]). This is usually done by sending Bayesian inference jobs to high-performance computing (HPC) clusters, making use of large numbers of CPU cores to obtain the desired parallelization gains. However, even for these implementations the computational bottleneck can be found in the calculation of the likelihood, in particular the generation of the waveform. Methods like RIFT can also benefit from fast waveform generation, though to a lesser extent, as they require a single step of waveform generation over a pre-defined grid [17].
@@
Next-generation GW detectors, namely Einstein Telescope and Cosmic Explorer, are expected to achieve detection rates of 100,000 BBH events per year [18, 19]. In order to process these events efficiently with Bayesian methods, the waveform generation process must be made faster, while still allowing for an accurate parameter estimation. In order to address this issue, applications of deep learning methods to steps of this process have been proposed. Approaches based on likelihood-free inference avoid live waveform generation altogether by replacing Bayesian parameter estimation with neural posterior estimation methods (though efficient waveform generators are still useful for the training process) [20, 21], while others strive for a speed-up through more efficient sampling using normalizing flows [22, 23].
In this paper, we instead tackle the problem through the lens of speeding up waveform generation. Existing literature has shown machine learning to be an effective method for waveform generation, in particular to provide reduced order models of existing approximants. Previous works have successfully demonstrated the generation of accurate waveforms through the mapping of physical parameters to the coefficients of a reduced order basis obtained through greedy algorithms [24, 25, 26, 27], an approach which has been reproduced for PCA bases [28, 29], as well as SVD bases  [30, 31, 32].
We expand on previous works by introducing a strategy for the training of fast and accurate neural surrogate models for numerical relativity data. By training a small neural network to predict the coefficients of a reduced PCA basis of approximant waveforms, and then fine-tuning on numerical relativity waveforms, allowing the network to dynamically adjust the basis, it is possible to obtain a model which is capable of generating millions of accurate waveforms on a consumer-level GPU in a matter of milliseconds. As a demonstration of this training strategy, we present  DANSur3dq8  (Dual-stage Artificial Neural Surrogate), a fast and accurate surrogate to the dominant (2,2) mode of numerical-relativity-produced GWs leveraging the GPU-based nature of modern machine learning. As a show of generalizability, we produce 4 versions of DANSur3dq8, pre-training on datasets obtained from 4 different GW approximants.
@@
The paper is organized as follows: in Section II, we introduce the datasets used in this work, including the different approximants and numerical relativity waveforms. Section III describes our methodology, detailing the dimensionality reduction approach, neural network architecture used for training and the training workflow. The results of our model, including accuracy, speed, and parameter estimation capabilities, are presented in Section IV. Finally, we summarize our findings and discuss potential future directions in Section V.
The code developed to create and train the models used in this work can be found at https://github.com/osvaldogramaxo/DANSur_22/.
@@ II Datasets
In the development of deep learning methods, data availability often poses a significant constraint. This proves particularly problematic in the field of numerical relativity, where the high costs associated with simulations result in a scarcity of data. For instance, by applying constraints of an effective precessing spin $\chi_{p}<0.001.$ and a mass ratio $q<=8$ to the SXS BBH data [10], we obtain only 376376376 samples of numerical relativity simulations. Given that deep learning frameworks typicaly need very large datasets in order to properly interpolate between the points of the input and output parameter spaces, current numerical catalogues are inadequate to fully explore them. To compensate for this, we propose to generate a substantial GW dataset using a existing approximants to train a model which can capture the fundamental characteristics of BBH GW signals, and subsequently fine-tune this model using numerical relativity data. In our study we will use four approximants: NRHybSur3dq8, a best-in-class surrogate model built using an empirical interpolant [33], SEOBNRv5HM, an approximant making use of the effective-one-body (EOB) formalism in order to generate GW waveforms [34], SEOBNRv4HM_PA, an EOB approximant accelerated by using a post-adiabatic approach to evolve the system dynamics [35], and IMRPhenomTHM, an approximant making use of phenomenological fits to generate waveforms in the time domain [36]. Data for the latter two is generated through their lalsimulation implementation [37], data for NRHybSur3dq8 is generated using the  gwsurrogate package for python [38], and data for SEOBNRv5HM is generated with the pyseobnr package [39].
II.1 Approximant datasets
@@
For each approximant, we generate 102,400 waveforms. We restrict ourselves to aligned-spin BBH mergers, in order to keep complexity manageable, with mass ratios sampled uniformly from the [1,8][1,8][ 1 , 8 ] interval, while the dimensionless component spins $\chi_{1},\chi_{2}$ are sampled uniformly from the [0.8,0.8][-0.8,0.8][ - 0.8 , 0.8 ] interval. We note that while uniform sampling does not necessarily reflect population-informed priors, it avoids biasing towards any particular parameter combination. The waveforms are generated in simulation units, with amplitude and time scaled by the sum of the Christodoulou masses of both black holes, with a corresponding sample rate of $2~\mathrm{M}_{\odot}^{-1}$. The merger time is estimated as the time which maximizes the amplitude of the quadrupolar $(l,|m|)=(2,2)$ mode, and the waveforms are trimmed to a duration of $4096~\mathrm{M}_{\odot}$, lining up the merger $100~\mathrm{M}_{\odot}$ before the end of the window. We further simplify our problem by setting the initial phase of the (2,2) mode in all the waveforms to zero. The (2,2) mode of the waveform is then saved to an HDF5 file, along with the physical parameters for the generation.
@@ II.2 Numerical Relativity waveforms
For setting up the numerical relativity dataset, we make use of the publicly available BBH simulation data from the SXS collaboration [10], consisting of 2018 simulations. The (2,2) waveforms are extracted using the SXS python package [40]. To match the data we produced with the approximant, we must do a number of checks. First, we take into account only the waveforms which have a duration of at least $4096~\mathrm{M}_{\odot}$. Then, we select the non-precessing waveforms by filtering effective precessing spins $\chi_{p}<0.001$ [41], with mass ratios $q\leq 8$, so that we are in line with the approximant data described previously. We also make sure memory effects are not present. This process leaves us with only 376 waveforms. These waveforms are interpolated using cubic splines to the same sample rate as the approximant dataset, and trimmed to $4096~\mathrm{M}_{\odot}$.
@@ III.1 Dimensionality reduction
When dealing with numerical data, gravitational waveform modes are essentially arrays of $M$-length complex numbers. Formally, these are points in $\mathbb{C}^{M}$, which can be reframed as points in $\mathbb{R}^{N}$ with $N=2M$. However, the high correlation between adjacent points in the waveform, as well as restrictions to the shape of the waveforms (that is, wavelike signals with a chirp and ringdown-like features) means that gravitational waveforms occupy a very restricted subspace of $\mathbb{R}^{N}$. On the other hand, neural networks with $N$ outputs will output to some generic subspace of $\mathbb{R}^{N}$, the size of which will depend on the number of layers and neurons per layer (that is, its expressivity). Therefore, we need the network to learn not only the shape of the subspace, as well as how this space is connected. Furthermore, the average Euclidean distance between two points in a cube of side $l$ in $\mathbb{R}^{N}$ scales as $O(\sqrt{N})$ [42] .
@@
This means that, in general, high dimensional data will be more sparse than lower dimensional data. As such, a neural network will have to interpolate over longer distances in such a case, which may lead to problems with convergence and generalisation [43]. Therefore, the projection of $N$-dimensional data to some lower-dimensional space should make the data more dense and thus facilitate the ease of convergence of neural networks training on that data. This is the concept of dimensionality reduction, and it is widely used across applications of machine learning [44, 45, 46]. There are many methods for dimensionality reduction, both linear and non-linear. Our use case involves training a model in a lower-dimensional representation space of gravitational waves, and then performing an inverse transform from that space back to the typical waveform representation. This requires that any method we choose must have an accurate inverse transform. While linear methods typically have an easily calculable inverse, non-linear methods very often do not have a tractable inverse, and they must be fitted to the data in a similar way as the forward transform, such as in autoencoders or UMAP [47, 48]. In this work, we will focus on principal component analysis (PCA), one of the most successful linear dimensionality reduction methods [49, 50]. It is worth noting that SVD works similarly to PCA and as such it would be expected to yield similar results. Similarly, other dimensionality reduction methods such as the aforementioned greedy basis methods can be used in a similar way. While a thorough comparison of dimensionality reduction methods for the creation of surrogate models is lacking in literature, the issue has been discussed in theory at length in [51], while a comparison of SVD and greedy basis methods can be found in appendix B of [52].
Figure 2: Mean mismatch values as a function of the number of principal components kept for the amplitude and phase. The arrows show the gradient of the mismatch values, and the yellow circle shows the chosen number of components to build the PCA basis.
@@
PCA works by calculating a linear transform $\mathcal{L}:X~\rightarrow(X-\mu)V$, where $X$ is some input $N\times M$ matrix (interpreted in our case as being a collection of $N$ waveforms of length $M$), $\mu$ is the average value at each timestep, and $V$ is the $m\times n$ matrix of eigenvectors of the covariance matrix $\Sigma=\frac{1}{n}(X-\mu)^{T}(X-\mu)$. The crucial point here is that the centering of $X$ means that the eigenvalues of $\Sigma$ represent the explained variance of the corresponding eigenvector. As such, if we consider only the eigenvalue-sorted subset of the top-valued $k$ eigenvectors, $V_{k}$ (an $m\times k$ matrix), we can project the data onto the reduced basis of dimension $k$ through the PCA transform:
$X^{PCA}_{K}=(X-\mu)V_{k}.$ (1)
We can then approximately reconstruct the original dataset through the inverse PCA transform
$X^{\prime}=X^{PCA}_{K}V^{T}_{k}+\mu\,.$ (2)
@@
There will be some error $E(X,X^{\prime})$ associated to this reconstruction, which will depend on the amount of components truncated (that is, $E(X,X^{\prime})\rightarrow 0$ as $k~\rightarrow n$).
@@
In our case, we will separate the complex waveform arrays $h=h_{+}-ih_{\times}=Ae^{i\phi}$ into 2-channel arrays, separating the amplitude $\mathcal{A}$ and (unwrapped) phase $\varphi$ quantities, as represented in Figure 1. This allows for a simpler, numerically non-oscillating representation, and also allows us to perform PCA separately on each quantity. Therefore, we can tune the amount of precision required in the reconstruction of each quantity. In Figure 2 we show the average mismatch values (see definition below) and their gradients as a function of the number of principal components kept for the amplitude and phase, calculated for the  NRHybSur3dq8 dataset. As can be seen, the influence of the number of phase components slightly outweighs the amplitude, and after 19 amplitude components and 23 phase components, the gradients essentially vanish. Going forward, our chosen representation will thus keep 19 components for the amplitude and 23 components for the phase, thus reducing the 2048-dimensional complex representation of waveforms (which is equivalent to a 4096-dimensional representation) to a 42-dimensional one.